# Efficient Neuro-Symbolic Learning of Constraints & Objective

### v1

@@ 2.2. Approaches without a discrete model
Other related approaches drift away from the DFL setting as they do not explicitly model or assume the existence of an underlying optimisation model that could be learned. A neural architecture directly learns to predict good solutions $\mathbf{y}$ from $\omega$. These approaches can naturally not guarantee constraint satisfaction and, unless properly conditioned, need to be retrained if constraints need to evolve. This area of research has recently developed with the observed incapacity of autoregressive Large Language Models (LLMs) to learn how to solve problems that require thinking ahead, such as Sudoku (Ye et al., 2025) or Zebra-like problems (Smith, 1992; Lin et al., 2025; Zubic et al., 2025), as well simple path finding problems (Bachmann and Nagarajan, 2024). The seminal Recurrent Relational Network paper (Palm et al., 2018) uses a Graph Neural Net-like network to learn how to solve Sudoku. Later, a recurrent Transformer architecture (Yang et al., 2023) was trained to solve Sudoku and Visual Sudoku. Finally, as an alternative to autoregressive LLMs, a discrete Denoising Diffusion Probabilistic Model (DDPM) was recently proposed to solve the same tasks (Ye et al., 2025), with a 100% accuracy.
Stochastic GMs and deep learning have often been combined in machine learning and image processing in architectures that resemble ours, but with very different targets, such as semantic image segmentation (Liu et al., 2017) or semi-supervised labeling (Qu et al., 2019).
Overall, in the context of learning how to solve NP-hard discrete optimisation problems, our approach is the only one that requires no solver call at training while learning a model that can be solved with guarantees, and the only one that targets an objective which is non-linear in its decision variables while being able to learn an objective together with constraints. In its ability to learn constraints, it is closely related to the work of (Li et al., 2023b) that combines neural networks with the MaxSAT Z3 solver (De Moura and Bjørner, 2008). The main difference lies in the training procedure: we use an efficient convex loss while they rely on a “Difference of Convex” programming formulation.
@@ 3.1. Background on discrete Graphical Models
A discrete graphical model (GM) is a concise description of a joint function of many discrete variables as the combination of many simple functions. Depending on the nature of the output of the functions (Boolean or numerical), and how they are combined and described, GMs cover a large spectrum of AI NP-hard reasoning and optimization frameworks including Constraint Networks, Propositional Logic as well as their numerical additive variants Cost Function Networks and partial weighted MaxSAT (Cooper et al., 2020; Schiex et al., 1995). Following cost exponentiation and normalization, these numerical joint functions can describe joint probability distributions, as done in Markov Random Fields (MRFs). In this paper, we use Cost Function Networks for their ability to express both numerical and logical functions. We assume here that cost functions take their value in $\bar{\mathbb{R}}=\mathbb{R}\cup\{\infty\}$.
@@
Given a sequence $\mathbf{Y}=\{Y_{1},\ldots,Y_{n}\}$ of $n$ finite domain variables, a cost function network $\mathcal{M}$ is defined as a set of elementary cost functions. It defines a joint cost function, also denoted $\mathcal{M}(\cdot)=\sum_{F\in\mathcal{M}}F$, involving all variables in $\mathbf{Y}$. The optimization problem, known as the Weighted Constraint Satisfaction Problem (WCSP), is to find an assignment $\mathbf{y}$ that minimizes the joint function $\mathcal{M}(\mathbf{y})$. If $\mathcal{M}(\mathbf{y})<\infty$, $\mathbf{y}$ is called a solution (a model in propositional logic).
@@
Consider $\mathbf{Y}=\{Y_{1},Y_{2},Y_{3},Y_{4}\}$ with domains $\{0,1\}$ and $\mathbf{C}=\{Y_{1}\neq Y_{2},Y_{2}+Y_{3}>1,Y_{3}\neq Y_{4}\}$ represented in Figure 1. No constraint is redundant in $\mathbf{C}$, but in the context of the assignment $\{Y_{2}=1,Y_{3}=1\}$, the constraint $Y_{2}+Y_{3}>1$ becomes redundant w.r.t. $\mathbf{C}^{\prime}=\mathbf{C}\cup\{Y_{2}=1,Y_{3}=1\}$ . In the context of $\{Y_{1}=0\}$, $Y_{2}+Y_{3}>1$ becomes partially redundant, as it could equivalently be replaced by the weaker $Y_{2}=Y_{3}$. Due to redundancies, the observed values in a sample can create a context that makes some constraints redundant and, therefore, not learnable.
For $n$ variables, a strictly pairwise graphical model $\mathcal{M}$ ($\forall F\in\mathcal{M}$, $F$ involves exactly $2$ variables) can be described with n(n1)/2n(n-1)/2 elementary cost functions with tensors (matrices) of size at most $d^{2}$. We denote by $\mathcal{M}[i,j]$ the tensor describing the cost function between variables $Y_{i}$ and $Y_{j}$ in $\mathcal{M}$. In practice when $n$ is large, we can model a restricted number of cost functions (see Subsection 5.4 on protein design).
@@ 3.2. Problem statement
In this work, we assume that we observe samples $(\omega,\mathbf{y})$ of the values $\mathbf{y}$ of the variables $\mathbf{Y}$ as low-cost solutions of an underlying constrained optimization problem with parameters influenced by natural inputs $\omega$. From a data set $\mathbf{S}$ of pairs $(\omega,\mathbf{y})$, we want to learn a model $N$ (in our case, a neural network) that predicts a pairwise graphical model $\mathcal{M}=N(\omega)$ such that $\mathbf{y}\in\operatorname*{arg\,min}_{\mathbf{y}\in D^{\mathbf{Y}}}N(\omega)(\mathbf{y})$. This graphical model $\mathcal{M}=N(\omega)$ defines the last layer of our hybrid neural+graphical model architecture (see Figure 2).
Figure 2. Our hybrid learning architecture: natural inputs $\omega$ (left) feed a neural net NN in charge of predicting all pairwise cost functions $F_{ij}$ of the GM $\mathcal{M}=N(\omega)$. To learn NN, we back-propagate solutions $\mathbf{y}$ for $(\omega,\mathbf{y})\in\mathbf{S}$ through the E-PLL loss function. At inference, $N(\omega)$ can be directly fed to any GM solver, be it exact, based on a scalable relaxation or a (meta)-heuristic. This is illustrated here on 2 possible problems: a visual Sudoku problem (top) and a protein design problem (bottom).
@@
Assuming the data set $\mathbf{S}$ contains i.i.d. samples from an unknown probability distribution $P(\mathbf{Y}|\omega)$, a natural loss function for the GM $N(\omega)$ is the asymptotically-consistent negative logarithm of the probability $\textrm{NLL}(\mathbf{S})=-\log(\prod_{(\omega,\mathbf{y})\in\mathbf{S}}P^{N(\omega)}(\mathbf{Y}=\mathbf{y}))$ of the observed samples. This negative loglikelihood is, however, intractable because of the #P-hard normalizing constant needed to compute the probability. We instead rely on the tractable negative pseudo-loglikelihood (Besag, 1975) $\textrm{NPLL}(S)=\sum_{(\omega,y)\in S}-\log(\prod_{i}P^{N(\omega)}(Y_{i}=y_{i}|\mathbf{y}_{-i}))$. The NPLL works at the level of each variable $Y_{i}$, in the context of $\mathbf{y}_{-i}$, the assignment of all other variables. It requires only normalization over one variable $Y_{i}$, a computationally easy task. Generalized variants of the NPLL use subgroups of variables (Huang and Ogata, 2002) instead of single variables, at the cost of additional complexity. In Appendix B.1, we give a simple proof of the asymptotic consistency of the basic PLL (a folklore result), under a strict convexity assumption, while Appendix B.2 reframes the PLL as a Fenchel-Young loss.
@@
Due to the decomposability of the joint function $\mathcal{M}(\cdot)$, we have $P^{\mathcal{M}}(Y_{i}|\mathbf{y}_{-i})=\mathrm{softmax}(-m_{i}(Y_{i}))$ where $m_{i}(Y_{i})=\sum_{j\neq i}\mathcal{M}[ij](Y_{i},y_{j})$. In a message passing interpretation, $m_{i}(\cdot)\in\bar{\mathbb{R}}^{|D^{i}|}$ represents the sum of all messages received from neighbor variables $Y_{j}$ through the incident functions $\mathcal{M}[i,j]$, given $Y_{j}=y_{j}$. Computing the NPLL is in O(n(n1)d)O(n(n-1)d) per sample and epoch. It can easily be vectorized (computed independently on each variable).
The NPLL enables scalable training from natural inputs in the context of an underlying NP-hard GM optimization problem. However, the proofs of asymptotic consistency of the NPLL (Besag, 1975; Geman and Graffigne, 1986) relies on identifiability or strict convexity assumptions (see Appendix B.1) that do not hold in the context of constraints (zero probabilities) as redundant constraints can be added or not without influencing the joint function $\mathcal{M}(\cdot)$. Unsurprisingly, the NPLL is known to perform poorly in the presence of large costs (Montanari and Pereira, 2009; Huang and Ogata, 2002). Empirically, we observed that the resulting architecture completely fails at solving even the simplest symbolic Sudoku problem where $\omega$ contains the digits from the unsolved Sudoku grid and $\mathbf{y}$ is the corresponding solution.
@@ 4. The E-PLL
The issue with the NPLL lies in the dynamic of the stochastic gradient optimization: the early identification of some high costs under $\omega$ will prevent the increase of other significant costs which are redundant in the context of the observed $\mathbf{y}$, but not redundant in the global unconditioned problem.
Inspired by “dropout” in deep learning (Srivastava et al., 2014), we introduce the Emmental NPLL (E-PLL) as an alternative to the NPLL that can learn all constraints (infeasibilities) present in $S$.
@@
The idea of the E-PLL follows directly from the previous gradient analysis: to prevent a combination of incident functions with already-learned high cost from shrinking gradients, we mute a random fraction of the incident functions. It can be understood intuitively on the Sudoku example: the NPLL asks to predict one cell given the rest of the grid. Once all the row constraints are learnt, one missing cell will always be predicted correctly and the loss is zero. Yet for inference all the constraints are needed. To force the neural network to also learn column and square constraints, the E-PLL randomly masked some of the cells.
The E-PLL is used alongside an L1 regularization on the output of the learned network $N(\omega)$ to favour sparsity. This also makes the GM optimization problem easier to solve. The complete training loss is then:
@@ 4.1. Dealing with unobserved variables
The E-PLL is defined only over complete assignments. In the context of missing data, in every sample $(\omega,\mathbf{y})$, $\mathbf{y}$ may be a partial assignment. One example is the visual Sudoku, were each image of a digit in the grid is known to represent the value of a single variable. If the values of initial hints are observed in $\mathbf{y}$, they provide direct supervision and thus grounding information (Chang et al., 2020) when training digit recognition. Simultaneous learning of Sudoku rules and digit recognition requires to mask initial hints in the sequence, leading to unobserved variables. We tackle this Visual Sudoku problem in Subsection 5.2.
A usual strategy to deal with missing data is to rely on variants of the expectation-maximization (EM) algorithm (Qu et al., 2019), where expectations of sufficient statistics (Kholevo, 2001) are used in the likelihood maximisation algorithm. In our case, this would require computing expensive #-P hard marginals. We instead rely on a simpler, NP-hard, imputation procedure where the values of the missing variables are obtained by optimizing the joint function defined by $N(\omega)$, all variables observed in $\mathbf{y}$ being assigned to their values in $\mathbf{y}$. The resulting complete assignment is then used instead of the partial $\mathbf{y}$. During training, this strategy requires one NP-hard oracle call per sample with missing data. When the fraction of unobserved variables remains limited, the solved problems are simple, with just a few unassigned variables. In our experiments, we use an exact GM solver to obtain a single imputation. Heuristic or approximate solvers (Mladenović and Hansen, 1997; Durante et al., 2022) could be used instead, with multiple imputations (Little and Rubin, 2019) if necessary.
@@ 5. Experiments
    Learning the constraints of purely-symbolic logical puzzles: Sudoku and Futoshiki, with one or many solutions (Nandwani et al., 2021)
@@
    Learning the constraints of logical puzzles (Sudoku) with visual inputs, a task known as Visual Sudoku (Wang et al., 2019). We consider the ungrounded setting where the observed digit images in $\omega$ are missing in the corresponding $\mathbf{y}$.
@@
Reported training times have been measured using a Nvidia RTX-2070 Super GPU with 8GB of VRAM and a $4.2$ GHz AMD Ryzen 9 5900X CPU with 32 GB of RAM. During our measures, we noticed that the training accuracies were often slightly below those reported in the original papers. For all those with training times below 50 hours, we reran them twice to collect a total of 3 accuracies (including the paper result). For these, we report the min/max accuracies observed (when they differ). Our code is run with PyTorch 2.6 and PyToulbar2 0.0.0.4. We use the Adam optimizer with a weight decay of $10^{-4}$ and a learning rate of $10^{-3}$ (other parameters take default values). An L1 regularization with multiplier $\lambda=2.10^{-4}$ is applied on the cost matrices $N(\omega)[i,j]$. Code and data are available at https://github.com/mdefresne/emmental_pll/.

### v3

@@ 2.2. Approaches without a Discrete Model
Other related approaches drift away from the DFL setting as they do not explicitly model or assume the existence of an underlying optimisation model that could be learned. A neural architecture directly learns to predict good solutions $\mathbf{y}$ from $\omega$. These approaches can naturally not guarantee constraint satisfaction and, unless properly conditioned, need to be retrained if constraints need to evolve. This area of research has recently developed with the observed incapacity of autoregressive Large Language Models (LLMs) to learn how to solve problems that require thinking ahead, such as Sudoku (ye2025beyond, 75) or Zebra-like problems (smith1992solve, 63, 39, 76), as well simple path finding problems (DBLP:conf/icml/BachmannN24, 6). The seminal Recurrent Relational Network paper (RRN, 54) uses a Graph Neural Net-like network to learn how to solve Sudoku. Later, a recurrent Transformer architecture (Yang2023learning, 73) was trained to solve Sudoku and Visual Sudoku. Recently, as an alternative to autoregressive LLMs, a discrete Denoising Diffusion Probabilistic Model (DDPM) was proposed to solve the same tasks (ye2025beyond, 75), with a 100% accuracy and a related Hierarchical Reasoning model (wang2025hierarchical, 69) shown to be able to solve extremely challenging Sudoku grids with 55% accuracy.
Stochastic GMs and deep learning have often been combined in machine learning and image processing in architectures that resemble ours, but with very different targets, such as semantic image segmentation (liu2017deep, 42) or semi-supervised labeling (qu2019gmnn, 58).
Overall, in the context of learning how to solve NP-hard discrete optimisation problems, our approach is the only one that requires no solver call at training while learning a model that can be solved with guarantees, and the only one that targets an objective that is non-linear in its decision variables while being able to learn an objective together with constraints.
@@ 3.1. Background on Discrete Graphical Models
A discrete graphical model (GM) is a concise description of a joint function of many discrete variables as the combination of many simple functions. Depending on the nature of the output of the functions (Boolean or numerical), and how they are combined and described, GMs cover a large spectrum of AI NP-hard reasoning and optimisation frameworks, including Constraint Networks, Propositional Logic as well as their numerical additive variants Cost Function Networks and partial weighted MaxSAT (cooper2020graphical, 16, 61). Following cost exponentiation and normalization, these numerical joint functions can describe joint probability distributions, as done in Markov Random Fields (MRFs). In this paper, we use Cost Function Networks for their ability to express both numerical and logical functions. We assume here that cost functions take their value in $\bar{\mathbb{R}}=\mathbb{R}\cup\{\infty\}$.
@@
Given a sequence $\mathbf{Y}=\{Y_{1},\ldots,Y_{n}\}$ of $n$ finite domain variables, a cost function network $\mathcal{M}$ is defined as a set of elementary cost functions. It defines a joint cost function, also denoted $\mathcal{M}(\cdot)=\sum_{F\in\mathcal{M}}F$, involving all variables in $\mathbf{Y}$. The optimisation problem, known as the Weighted Constraint Satisfaction Problem (WCSP), is to find an assignment $\mathbf{y}$ that minimizes the joint function $\mathcal{M}(\mathbf{y})$. If $\mathcal{M}(\mathbf{y})<\infty$, $\mathbf{y}$ is called a solution (a model in propositional logic).
@@
Consider $\mathbf{Y}=\{Y_{1},Y_{2},Y_{3},Y_{4}\}$ with domains $\{0,1\}$ and $\mathbf{C}=\{Y_{1}\neq Y_{2},Y_{2}+Y_{3}>1,Y_{3}\neq Y_{4}\}$ represented in Figure 1. No constraint is redundant in $\mathbf{C}$, but in the context of the assignment $\{Y_{2}=1,Y_{3}=1\}$, the constraint $Y_{2}+Y_{3}>1$ becomes redundant w.r.t. $\mathbf{C}^{\prime}=\mathbf{C}\cup\{Y_{2}=1,Y_{3}=1\}$ . In the context of $\{Y_{1}=0\}$, $Y_{2}+Y_{3}>1$ becomes partially redundant, as it could equivalently be replaced by the weaker $Y_{2}=Y_{3}$. Due to redundancies, the observed values in a sample can create a context that makes some constraints redundant and, therefore, not learnable as its presence or absence does not change the problem semantics.
For $n$ variables, a strictly pairwise graphical model $\mathcal{M}$ ($\forall F\in\mathcal{M}$, $F$ involves exactly $2$ variables) can be described with n(n1)/2n(n-1)/2 elementary cost functions with tensors (matrices) of size at most $d^{2}$. We denote by $\mathcal{M}[i,j]$ the tensor describing the cost function between variables $Y_{i}$ and $Y_{j}$ in $\mathcal{M}$. In practice, when $n$ is large, we can model a restricted number of cost functions (see Subsection 5.4 on protein design).
@@ 3.2. Problem Statement
In this work, we assume that we observe samples $(\omega,\mathbf{y})$ of the values $\mathbf{y}$ of the variables $\mathbf{Y}$ as low-cost solutions of an underlying constrained optimisation problem with parameters influenced by natural inputs $\omega$. From a data set $\mathbf{S}$ of pairs $(\omega,\mathbf{y})$, we want to learn a model $N$ (in our case, a neural network) that predicts a pairwise graphical model $\mathcal{M}=N(\omega)$ such that $\mathbf{y}\in\operatorname*{arg\,min}_{\mathbf{y}\in D^{\mathbf{Y}}}N(\omega)(\mathbf{y})$. This graphical model $\mathcal{M}=N(\omega)$ defines the last layer of our hybrid neural+graphical model architecture (see Figure 2).
Figure 2. Our hybrid learning architecture: natural inputs $\omega$ (left) feed a neural net NN in charge of predicting all pairwise cost functions $F_{ij}$ of the GM $\mathcal{M}=N(\omega)$. To learn NN, we back-propagate solutions $\mathbf{y}$ for $(\omega,\mathbf{y})\in\mathbf{S}$ through the E-PLL loss function (no solver is used). At inference, $N(\omega)$ can be directly fed to any GM solver, be it exact, based on a scalable relaxation or a (meta)-heuristic. This is illustrated here on 2 possible problems: a visual Sudoku problem (top) and a protein design problem (bottom).
@@
Assuming the data set $\mathbf{S}$ contains i.i.d. samples from an unknown probability distribution $P(\mathbf{Y}|\omega)$, a natural loss function for the GM $N(\omega)$ is the asymptotically-consistent negative logarithm of the probability $\textrm{NLL}(\mathbf{S})=-\log(\prod_{(\omega,\mathbf{y})\in\mathbf{S}}P^{N(\omega)}(\mathbf{Y}=\mathbf{y}))$ of the observed samples. This negative loglikelihood is, however, intractable because of the #P-hard normalizing constant needed to compute the probability. We instead rely on the tractable negative pseudo-loglikelihood (besag1975statistical, 9) $\textrm{NPLL}(S)=\sum_{(\omega,y)\in S}-\log(\prod_{i}P^{N(\omega)}(Y_{i}=y_{i}|\mathbf{y}_{-i}))$. The NPLL works at the level of each variable $Y_{i}$, in the context of $\mathbf{y}_{-i}$, the assignment of all other variables. It requires only normalization over one variable $Y_{i}$, a computationally easy task. Generalized variants of the NPLL use subgroups of variables (huang2002generalized, 27) instead of single variables, at the cost of additional complexity. In Appendix B.1, we give a simple proof of the asymptotic consistency of the basic PLL (a folklore result), under a strict convexity assumption, while Appendix B.2 reframes the PLL as a Fenchel-Young loss.
@@
Due to the decomposability of the joint function $\mathcal{M}(\cdot)$, we have $P^{\mathcal{M}}(Y_{i}|\mathbf{y}_{-i})=\mathrm{softmax}(-m_{i}(Y_{i}))$ where $m_{i}(Y_{i})=\sum_{j\neq i}\mathcal{M}[i,j](Y_{i},y_{j})$. In a message passing interpretation, $m_{i}(\cdot)\in\bar{\mathbb{R}}^{|D^{i}|}$ represents the sum of all messages received from neighbor variables $Y_{j}$ through the incident functions $\mathcal{M}[i,j]$, given $Y_{j}=y_{j}$. Computing the NPLL is in O(n(n1)d)O(n(n-1)d) per sample and epoch. It can easily be vectorized (computed independently on each variable).
The NPLL enables scalable training from natural inputs in the context of an underlying NP-hard GM optimisation problem. However, the proofs of asymptotic consistency of the NPLL (besag1975statistical, 9, 25) rely on identifiability or strict convexity assumptions (see Appendix B.1) that do not hold in the context of constraints (zero probabilities) as redundant constraints can be added or not without influencing the joint function $\mathcal{M}(\cdot)$. Unsurprisingly, the NPLL is known to perform poorly in the presence of large costs (difficultGM, 49, 27). Empirically, we observed that the resulting architecture completely fails at solving even the simplest symbolic Sudoku problem where $\omega$ contains the digits from the unsolved Sudoku grid and $\mathbf{y}$ is the corresponding solution.
@@ 4. The E-PLL
The issue with the NPLL lies in the dynamic of the stochastic gradient optimisation: the early identification of some high costs under $\omega$ will prevent the increase of other significant costs which are redundant in the context of the observed $\mathbf{y}$, but not redundant in the global unconditioned problem.
Inspired by “dropout” in deep learning (srivastava2014dropout, 64), we introduce the Emmental NPLL (E-PLL) as an alternative to the NPLL that can learn all constraints (infeasibilities) present in $S$.
@@
The idea of the E-PLL follows directly from the previous gradient analysis: to prevent a combination of incident functions with already-learned high cost from shrinking gradients, we mute a random fraction of the incident functions (or equivalently, variables, as we learn pairwise functions). It can be understood intuitively on the Sudoku example: the NPLL asks to predict one cell given the rest of the grid. Once all the row constraints are learnt, one missing cell will always be predicted correctly and the loss is zero. Yet for inference all the constraints are needed. To force the neural network to also learn column and square constraints, the E-PLL randomly masked some of the cells.
The E-PLL is used alongside an L1 regularization on the output of the learned network $N(\omega)$ to favour sparsity. This also makes the GM optimisation problem easier to solve. The complete training loss is then:
@@ 4.1. Dealing with Unobserved Variables
The E-PLL is defined only over complete assignments. In the context of missing data, in every sample $(\omega,\mathbf{y})$, $\mathbf{y}$ may be a partial assignment. One example is the visual Sudoku, were each image of a digit in the grid is known to represent the value of a single variable. If the values of initial hints are observed in $\mathbf{y}$, they provide direct supervision and thus grounding information (assessingSATNet, 15) when training digit recognition. Simultaneous learning of Sudoku rules and digit recognition requires to mask initial hints in the sequence, leading to unobserved variables. We tackle this Visual Sudoku problem in Subsection 5.2.
A usual strategy to deal with missing data is to rely on variants of the expectation-maximization (EM) algorithm (qu2019gmnn, 58), where expectations of sufficient statistics (kholevo2001sufficient, 31) are used in the likelihood maximisation algorithm. In our case, this would require computing expensive #P-hard marginals. We instead rely on a simpler, NP-hard, imputation procedure where the values of the missing variables are obtained by optimizing the joint function defined by $N(\omega)$, all variables observed in $\mathbf{y}$ being assigned to their values in $\mathbf{y}$. The resulting complete assignment is then used instead of the partial $\mathbf{y}$. During training, this strategy requires one NP-hard oracle call per sample with missing data: training scalability is, in principle, lost. This brings our approach closer to the Nesy-programming (li2023neuro, 36) approach, which allows for latent variables and needs to solve a (generally NP-hard) “Difference of Convex” problem.
When the fraction of unobserved variables remains limited, the solved problems are simple, with just a few unassigned variables. In our experiments, we use an exact GM solver to obtain a single imputation. For increased scalability, heuristic or approximate solvers (mladenovic1997vns, 48, 22) could be used instead, with multiple imputations (little2019statistical, 40) if necessary.
@@ 5. Experiments
    Learning the constraints of purely-symbolic logical puzzles: Sudoku and Futoshiki, with one or many solutions (1ofMany, 51)
@@
    Learning the constraints of logical puzzles (Sudoku) with visual inputs, a task known as Visual Sudoku (satnet, 70). We consider the ungrounded setting where the observed digit images in $\omega$ are missing in the corresponding $\mathbf{y}$.
@@
Reported training times have been measured using a Nvidia RTX-2070 Super GPU with 8GB of VRAM and a $4.2$ GHz AMD Ryzen 9 5900X CPU with 32 GB of RAM. During our measures, we noticed that the training accuracies were often slightly below those reported in the original papers. For all those that could run on our hardware with training times below 50 hours, we reran them twice to collect a total of 3 accuracies (including the paper result). For these, we report the min/max accuracies observed (when they differ). Our code is run with PyTorch 2.6 and PyToulbar2 0.0.0.4. We use the Adam optimizer with a weight decay of $10^{-4}$ and a learning rate of $10^{-3}$ (other parameters take default values). An L1 regularization with multiplier $\lambda=2.10^{-4}$ is applied on the cost matrices $N(\omega)[i,j]$. Code and data are available at https://github.com/mdefresne/emmental_pll/.